On August 1 of the current year, Kates employer Shop Inc. provided him a vehicle. For Shop Inc. to purchase the vehicle it would cost $45,000 including all applicable taxes. Whereas leasing the vehicle would cost Shop Inc. $750 a month. In either case, Shop Inc. would pay all operating costs for the vehicle, which is expected to be $2,250 yearly. Kates estimates that he will drive 2,100 monthly, of which 300 km will be for employment.

a) Determine the amount to be included in Kate's employment inform for the current year if i)

Shop Inc. purchases the car and ii) if Shop Inc. leases the car.

b) Kate drives the car for 16,000 km for employment purposes and 6,000 for personal use in the following year. Determine the amount to be included in his personal taxes assuming Shop Inc. leases the vehicle.
Answer #1

Ans. a). Total Yearly Cost when vehicle is purchased= $45000 &

Total Yearly Cost when vehicle is taken on lease= $750*12+2250= $11100

Total Run KM Yearly= 2100*12=25200 KMs

Total Run KM for Employment= 300*12=3600 KMs

Total Run KM for Private Purpose= 21600 KMs

The amount to be included in Employment inform when car is purchased; $45000*21600/25200= $38571.43

The amount to be included in Employment inform when car is on lease=$11100*21600/25200=$9514.29

Ans. b) In next year, total drive by Kate is 16000+6000=22000 KMs

Amount to be included in personal tax (when car is on lease) will be $11100*6000/22000= $3027.27